Solve the following system of equations by using Cramer's Rule: "-2x+3y-z=1," x+2y-z=4, "-2x-y+z=-3"
Answer by Alan3354(69443)

Solve the following system of equations by using Cramer's Rule: "-2x+3y-z=1," x+2y-z=4, "-2x-y+z=-3"
====================================
What is Cramer's Rule?
There's also Gauss, Gauss-Jordan, some others.
Why worry about what method is used?
I wrote an Excel sheet to do these. There's a site online that will do a 10 by 10 matrix, so I wasted my time.
Live and learn.
--------------------------
PS Is there some reason you can't do this assignment?